OpenStudy (anonymous):

(2 4/5) (1 1/4) perform the indicated operation.

OpenStudy (anonymous):

24*11/5*4 264/24 =11.

OpenStudy (anonymous):

no. (2 & 4/5) (1 & 1/4)

OpenStudy (anonymous):

Are these mixed numbers? (:

OpenStudy (anonymous):

than

OpenStudy (anonymous):

2 &4/5 =14/4 1 & 1/4 =5/4 14*5/4*4 = 70/16 35/8